Hello, need an explanation on how to find the x intercepts, and the vertex of this equation y is less than or equal to -x^2+3x+2. I do not know how to factor yet, so if we could do it by guessing two points and checking by plugging them into the equation and finding some easy to graph points. Thanks soOoOoOo much!
Answer by stanbon(57400)

need an explanation on how to find the x intercepts, and the vertex of this equation y is less than or equal to -x^2+3x+2.
--------------
a = -1 ; b = 3 ; c = 2
----
Vertex occurs at x = -b/(2a) = -3/(-2) = 3/2
---
Vertex y-value: f(3/2) = -(3/2)^2+3(3/2)+2 = -9/4+(9/2)+2 = 17/4
----
Vertex:(3/2 , 17/4)
--------------------------
x-intercepts:
solve -x^2+3x+2 = 0
x = [-3 +- sqrt(9-4*-1*2)]/(-2)
= [-3 +- sqrt(17)]/(-2)
Those are the x-intercepts
================================
Cheers,
Stan H.
============